OpenStudy (anonymous):

for \(f(x)=log_3(x-2)\) find \(f^{-1}(x)\)

OpenStudy (anonymous):

\[\log_3(y-2)=x\] solve for \(y\) in two steps

OpenStudy (anonymous):

woah woah woah....what?

OpenStudy (anonymous):

1) rewrite in equivalent exponential form 2) add 2 to both sides

OpenStudy (anonymous):

where did you get \(\log_3(y-2)=x\)?

OpenStudy (anonymous):

this is the way you are usually taught to find an inverse function right? if say \(f(x)=2x+3\) then to find \(f^{-1}(x)\) you write \[x=2y+3\] and solve for \(y\)

OpenStudy (anonymous):

ohhh...k i get it...gimme a sec

OpenStudy (anonymous):

i just did the same thing

OpenStudy (anonymous):

so it would be \(3^x=y-2\)\[3^x+2=y\]\[f^{-1}=3^x+2\]

OpenStudy (anonymous):

thanks! @satellite73 :)

OpenStudy (anonymous):

yw